# Understanding and Mitigating Membership Inference Risks of Neural Ordinary Differential Equations

### v1

@@ 5.4. NSDE as a Drop-in Replacement Module
All our evaluations so far assume that a defender trains NSDE models from scratch. This approach often requires more computational resources than expected, and could potentially face pushback, especially when well-performing pre-trained models are available. To address this issue, we present an efficient replace-then-finetune strategy. Instead of training NSDEs from scratch, we propose replacing the last few layers of an existing pre-trained model with an NSDE block and then finetune. This approach reduces computational demands by re-using the majority of pre-trained parameters, as it requires fewer training iterations, while also benefiting from NSDEs’ privacy advantages (based on DP’s composition theorem). It also potentially improves model utility, as fine-tuning leverages the features (i.e., the latent representations/activations obtained at the penultimate layer), a strategy demonstrated to be effective in prior work on training models with DP-SGD (Tramer and Boneh, 2021).
@@
Figure 7. Effectiveness of replace-and-finetune strategy. We replace the last residual block of ResNet14 with a NSDE block and then finetune the adapted model on CIFAR-10.
@@
Impact on privacy. Figure 7 demonstrates the effectiveness of the replace-then-finetune strategy. When the last group of the final residual block is replaced with an NSDE block, membership risks are reduced by more than an order of magnitude, particularly in low FPR regions, compared to ResNet-14. However, when compared to the NSDE models trained from scratch, membership risks are higher. This does not imply a difference in the privacy guarantees ($\varepsilon$) between the two models. Their guarantees remain the same due to the post-processing property. [^4] Note that the last NSDE block is randomly initialized and fine-tuned. It is likely that LiRA hits the worst-case privacy leakage against the replace-then-finetuned models, as indicated by the bound shown in Figure 5.
@@
Impact on model utility. We also note that the test accuracy of the replace-then-finetuned models is 84.0%, slightly decreased from 86.0% of ResNet-14. But the reduction is comparatively smaller than that we observe in NSDEs trained from scratch, which achieves a test accuracy of 81.9%. We also note that this approach potentially reduces the computational demands of NSDE training. Notably, we observe that test accuracy stabilizes after 150 epochs, which is half the total number of epochs we use for training NSDEs. As shown in our theoretical analysis in Sec 4.1, a reduction in the number of training iterations ($K$) also decreases the total privacy leakage of the trained models. Our results demonstrate the effectiveness of the replace-then-finetune strategy in reducing membership risks while achieving improved utility and lowering computational demands.
6. Related Work
@@
Privacy attacks on machine learning. Membership inference attacks, due to its simplest formulation—where an adversary knows a particular instance and then only aims to infer its membership, represent the worst-case privacy leakage and have been widely studied as a privacy auditing mechanism. But there have been other types of privacy attacks studied in machine learning.
@@
Fredrikson et al. (2014) introduced the concept of model inversion, which are reconstruction attacks designed to reverse-engineer training samples or reconstructing sensitive information in the training data not exposed to public. Initial work (Fredrikson et al., 2014) focused on attacks targeting simple linear models, while subsequent work (Fredrikson et al., 2015; Wu et al., 2016) focuses on non-linear models such as decision trees and/or neural networks. In particular, against neural networks, adversaries iteratively update instances by computing gradients, gradually refining them until they resemble the original training data.
@@
If a model inversion adversary focuses on inferring the value of unknown sensitive attributes, such as gender or race, of particular training records, it is referred to as attribute inference attacks (Yeom et al., 2018). A typical adversary exploits the query access to the target model to learn sensitive attributes in the (partial) training data. Diverse attack methodologies (Yeom et al., 2018; Mehnaz et al., 2022; Fredrikson et al., 2014) have been proposed, and attribute inference has been exploited across different domains, such as social media (Zheleva and Getoor, 2009; Chaabane et al., 2012; Gong et al., 2014; Gong and Liu, 2016; Jia et al., 2017; Gong and Liu, 2018) or recommendation systems (Weinsberg et al., 2012; Wu et al., 2020).
@@
The recent emergence of advanced neural networks, such as language models, diffusion models, and differential equation-based neural networks like NODEs, necessitates the adaptation and evaluation of existing privacy attacks on these models. The most active area of study focuses on language models and diffusion models. Prior work (Carlini et al., 2019, 2021, 2023) extended these privacy attacks to these models and demonstrated their vulnerability to membership inference attacks or model inversion, also referred to as data extraction attacks.
No prior work has yet studied the vulnerability of emerging differential equation-based neural networks, such as NODEs and NSDEs, to privacy attacks. Our work addresses this knowledge gap by presenting the first study on understanding and mitigating the membership risks associated with NODEs.
@@
Security and privacy of NODEs and NSDEs. In the literature, security and privacy aspects of NODEs and NSDEs have mainly focused on countering adversarial examples (Goodfellow et al., 2015) and learning stable, robust dynamics. Time-invariant steady neuarl ODEs (TisODEs) (Hanshu et al., 2020) empirically demonstrated that NODEs are more resilient to random Gaussian noise and to adversarial examples than traditional convolutional neural networks. TisODEs proposed a steady-state regularizer that minimizes the deviation of the hidden states ${}h(t)$ after a terminal time, $T$ (i.e., steady-state after $T$), leading to a stable dynamics. In (Kang et al., 2021), a method to guide trajectories toward Lyapunov-stable equilibrium points, resulting in NODEs extracting hidden features around these stable points. This work empirically shows that the stability leads to improved robustness against adversarial examples. The same approach has been further extended to a non-autonomous dynamics systems, where trajectories are guided to asymptotically stable equilibrium point (Li et al., 2022). In (Zeqiri et al., 2023), the investigation on the robustness has been extended to high-dimensional NODEs, presenting graph-based tools for robustness verification.
@@
Other techniques to enhance NODEs robustness include skew-symmetric ODE blocks (Huang et al., 2022) and integrating Guassian process (Anumasa and Srijith, 2021) both providing empirical improvements in robustness. While not explicitly robustness-focused, some efforts contribute similarly—regularizing NODE dynamics. Kinetic and Jacobian regularizers (Finlay et al., 2020) and high-order derivatives regularizers (Kelly et al., 2020) have been introduced and shown effective in regularizing the dynamics to be smooth. A method proposed in STEER (Ghosh et al., 2020) mitigate the learning of stiff ODEs by randomizing terminal times during training.
@@
There are only handful of papers that have investigated security and privacy aspects of NSDEs. Some of those work include (Jha et al., 2021), which showed that the attributions computed over NSDEs are less noisy, visually sharper, and quantitatively more robust than those computed on the deterministic NODEs. Liu et al. (2020) demonstrated that NSDEs improve robustness against adversarial examples. In general context, NSDEs have been widely adopted as a tool to estimate uncertainty (Kong et al., 2020; Kidger et al., 2021b), as a part of generative models (Kidger et al., 2021a), and physics simulation with probablistic state transitions (Boral et al., 2024).
7. Conclusion
@@
Our work studies the privacy implications of emerging neural networks, NODEs, using membership inference attacks as a vehicle. We are the first work that presents a comprehensive analysis of their membership risks. Our major findings are: (1) NODEs are twice as resilient to existing membership inference attacks compared to conventional feedforward models like ResNets, while achieving comparable accuracy. (2) This privacy benefit is not due to reduced model overfitting. NODEs exhibit a high generalization gap, and we attribute this benefit to their approach of modeling data as a system of ODEs. (3) In addition to that, we show that advanced, non-stochastic variants of NODEs offer reduced overfitting, resulting in an improved accuracy with a reduction in membership risks.
@@
Our analysis does not imply that NODEs are free from membership risks. We study an approach to further reduce the risks: NSDEs, a stochastic adaptation of NODEs that incorporates an additional diffusion term in system modeling. Most importantly, we formally show that the diffusion term works as a differentially-private (DP) mechanism, thereby making NSDEs as DP-leaners. Our theoretical analysis shows that the privacy guarantee ($\varepsilon$) offered by NSDEs is equivalent to that of DP-SGD when using identical mini-batch training configurations. We demonstrate empirically that NSDEs are more effective at reducing membership risks compared to existing defenses with no privacy guarantee. Unlike DP-SGD, they achieve this mitigation with a smaller impact on model utilty.
@@
Moreover, we present an effective strategy, replace-then-finetune, for using NSDEs as a drop-in replacement module to enhance privacy. We replace the last few layers of a pre-trained conventional feedforward model with an NSDE block and finetune the adapted model for a few epochs. We demonstrate that this strategy reduces the membership risks of the original pre-trained model by an order of magnitude, while achieving model utility greater than that of DP-SGD, with only half the computational costs for training.
@@
Our work demonstrates that systems learned by neural networks can be effectively modeled with privacy. Our results imply that advanced neural networks employing diffusion mechanisms, such as stable diffusion models, may, with careful modeling, operate as inherently private models while delivering enhanced performance. As future work, investigating this private system may offer valuable insights into how DP influences what neural networks learn. We also hope that our work inspires the adoption of the replace-then-finetune strategy as an effective approach to enhancing privacy.
Acknowledgements
@@
Appendix A Model Architectures
We adopt architectures examined in prior work (Oganesyan et al., 2020).
ResNet-14
$\begin{split}[\text{Conv2d}\rightarrow\text{BN}&\rightarrow\text{ResBlock}\times 3\rightarrow\text{BN}\rightarrow\text{Conv2d}\\ &\rightarrow\text{ResBlock}\times 3\rightarrow\text{BN}\rightarrow\text{Conv2d}\\ &\rightarrow\text{ResBlock}\times 3\\ &\rightarrow\text{AdaptiveAvgPool2d}\rightarrow\text{Linear}]\end{split},$
where ResBlock is composed as
@@ Appendix A Model Architectures
ODENet-16-32-64
$\begin{split}[\text{Conv2d}\rightarrow\text{BN}&\rightarrow\text{ODEBlock}\rightarrow\text{BN}\rightarrow\text{Conv2d}\\ &\rightarrow\text{ODEBlock}\rightarrow\text{BN}\rightarrow\text{Conv2d}\\ &\rightarrow\text{ODEBlock}\\ &\rightarrow\text{AdaptiveAvgPool2d}\rightarrow\text{Linear}]\end{split},$
where ODEBlock is
@@
ConcatConv2d augments $x$ by concatenating an additional feature channel derived from $t$.
The ResNet/ODENet model architecture consists of convolutional layers interspersed with batch normalization, integration blocks, downsampling, and fully connected layers.
The integration blocks are implemented as ResBlocks for ResNet and ODEBlocks for ODENet and SDENet. Each ResBlock or ODEBlock processes its inputs through a structured sequence of layers, adapting dynamically to varying input dimensions. Generally, a ResBlock begins with a Batch Normalization (BN) layer that normalizes the inputs, followed by a convolutional (Conv2d) layer. This convolutional (Conv2d) layer is configured to match the input dimension while maintaining the channel size, typically using a kernel size of 3×3, a stride of 1, and padding of 1 to preserve spatial dimensions. The convolution’s output undergoes another normalization via a second BN layer before being passed through an additional convolutional (Conv2d) layer with identical specifications to preserve feature depth and spatial integrity. Conversely, each ODEBlock incorporates time dynamics into the feature processing. It utilizes ConcatConv2d layers, which extend the input features by integrating a time channel, thus increasing the input channels by one. The first ConcatConv2d layer, therefore, takes in one additional channel compared to the feature channels—modifying the input dimension based on the incoming feature map size. This layer outputs feature maps that maintain the original dimensions, while the subsequent ConcatConv2d layer continues to process these features under the same dimensional constraints.
@@
Both models begin with a convolutional layer (Conv2d) that processes a 3-channel input image to produce 32 feature maps using a 3×3 filter with a stride of 2 and padding of 1. The resulting feature maps are then normalized through a Batch Normalization (BN) layer. Following the initial convolutional layer, the features pass through a group of either three ResBlocks or a single ODEBlock, depending on the configuration. Subsequently, the network transitions through a Downsample layer, which effectively reduces the spatial dimensions while adjust the channel depth, preparing the features for subsequent layers. For instance, after the first group of blocks, a downsample layer increases the channel depth from 32 to 64 while reducing the spatial resolution, preparing the data for the next set of deeper processing layers. Then the feature maps are processed through additional groups of three ResBlocks or an ODEBlock, followed by another downsample layers to transition the channel depth from 64 to 128. After the third group of three ResBlocks or an ODEBlock, the network utilizes an adaptive average pooling layer to condenses the feature maps into a single dimensional vector per feature map. This vector is then flattened and passed through a linear layer which acts as the classifier of the network, outputting the final predictions across the categories.
@@
Table 3. Membership inference risks of ODENet-64 models. We evaluate ODENet-64 models against five existing attacks on CIFAR-10 and CIFAR-100. We report the risks based on four different metrics employed in the prior work.
@@
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 ODENet-64 Yeom et al. (Yeom et al., 2018) 94.14% 84.36% 70.93% 51.85% 0.00% 0.03% 0.28% 1.17% 0.536 0.659 53.57% 66.09% Shokri et al. (Shokri et al., 2017) 0.08% 0.09% 0.75% 0.85% 0.473 0.480 50.03% 50.01% Song and Mittal (Song et al., 2021) 0.01% 0.13% 0.53% 1.58% 0.511 0.497 51.17% 51.63% Watson et al. (Watson et al., 2022) 0.00% 0.03% 0.28% 1.17% 0.536 0.659 53.57% 66.09% Carlini et al. (Carlini et al., 2022) 0.86% 1.72% 4.17% 7.70% 0.639 0.741 59.57% 69.49%
Appendix B Differential Privacy of Gaussian Mechanism

### v2

@@ 5.4 NSDE as a Drop-in Replacement Module
All our evaluations so far assume that a defender trains NSDE models from scratch. This approach often requires more computational resources than expected, and could potentially face pushback, especially when well-performing pre-trained models are available. To address this issue, we present an efficient replace-then-finetune strategy. Instead of training NSDEs from scratch, we propose replacing the last few layers of an existing pre-trained model with an NSDE block and then finetune. This approach reduces computational demands by re-using the majority of pre-trained parameters, as it requires fewer training iterations, while also benefiting from NSDEs’ privacy advantages (based on DP’s composition theorem). It also potentially improves model utility, as fine-tuning leverages the features (i.e., the latent representations/activations obtained at the penultimate layer), a strategy demonstrated to be effective in prior work on training models with DP-SGD [50]. In the following, we present results on CIFAR-10. We refer the readers to Appendix for results on other datasets.
@@
Figure 7: Effectiveness of replace-and-finetune strategy. We replace the last residual block of ResNet14 with a NSDE block and then finetune the adapted model on CIFAR-10.
@@
Impact on privacy. Figure 7 demonstrates the effectiveness of the replace-then-finetune strategy. When the last group of the final residual block is replaced with an NSDE block, membership risks are reduced by more than an order of magnitude, particularly in low FPR regions, compared to ResNet-14. However, when compared to the NSDE models trained from scratch, membership risks are higher. This does not imply a difference in the privacy guarantees ($\varepsilon$) between the two models. Their guarantees remain the same due to the post-processing property. [^4] Note that the last NSDE block is randomly initialized and fine-tuned. LiRA may hit the worst-case privacy leakage against the replace-then-finetuned models, as indicated by the bound shown in Figure 5.
@@
Impact on model utility. We also observe that the test accuracy of the replace-then-finetuned models is 84.0%, a modest decrease from the 86.0% achieved by ResNet-14. However, this reduction is comparatively smaller than that observed in NSDEs trained from scratch, which achieves a test accuracy of 81.9%. We emphasize that this approach also offers potentially reductions in the computational demands of NSDE training. Notably, we observe that test accuracy stabilizes after 150 epochs, which is half the total number of epochs we use for training NSDEs. As shown in our theoretical analysis in §4.1, a reduction in the number of training iterations ($K$) also decreases the total privacy leakage of the trained models. Our results demonstrate the effectiveness of the replace-then-finetune strategy in reducing membership inference risks while maintaining high utility and reducing computational overhead.
6 Related Work
@@
Privacy attacks on machine learning. Membership inference attacks—where an adversary, given a particular instance, attempts to infer whether it was part of a model’s training data—represent a worst-case form of privacy leakage and have been widely studied as a key mechanism for privacy auditing in machine learning. However, other types of privacy attacks have also been explored in the literature.
@@
Fredrikson et al. [51] introduced the concept of model inversion, a class of reconstruction attacks aimed at reverse-engineering training samples or recovering sensitive information in the training data that has not exposed to public. Initial work [51] focused on attacks targeting linear models, while subsequent work [52, 53] extended these attacks to non-linear models, including decision trees and neural networks. In particular, these attacks against neural networks typically involve iterative gradient-based optimization to refine inputs until they resemble the original training data.
@@
When a model inversion is used to infer sensitive attributes—such as gender or race—of specific training instances, it is referred to as attribute inference attacks [3]. These attacks exploit query access to a target model to extract unknown attributes from partially observed training records. Various attack methodologies [3, 54, 51] have been proposed, and attribute inference has been exploited across domains, such as social media [55, 56, 57, 58, 59, 60] or recommender systems [61, 62].
@@
The recent emergence of advanced neural networks, such as language models, diffusion models, and differential equation-based networks like NODEs, necessitates the adaptation and evaluation of existing privacy attacks on these models. While much attention has been given to language models and diffusion models, with prior studies [63, 64, 65] demonstrating their susceptibility to membership inference and model inversion (or data extraction) attacks, no prior work has systematically studied the privacy risks of differential equation-based neural networks such as NODEs and NSDEs. Our work addresses this gap by presenting the first study on understanding and mitigating the membership risks associated with differential equation-based models.
@@
Robustness of differential equation-based networks. In the existing literature, the security and privacy aspects of NODEs and NSDEs have mainly focused on countering adversarial examples [66] and ensuring the stability of learned dynamics. Time-invariant steady NODEs (TisODEs) [67], for instance, empirically demonstrated that NODEs are more resilient to random Gaussian noise and adversarial examples than traditional convolutional neural networks. TisODEs introduced a steady-state regularizer that minimizes the deviation of hidden states ${}h(t)$ beyond a terminal time $T$ (i.e., steady-state after $T$), effectively encouraging the network to learn stable dynamics. Similarly, Kang et al. [68] proposed guiding trajectory evolution toward Lyapunov-stable equilibrium points, enabling NODEs to extract features around stable attractors. This approach led to improved robustness against adversarial examples. The idea has been further extended to non-autonomous dynamical systems, guiding trajectories to asymptotically stable equilibrium points [69]. In [70], robustness analysis has also been extended to high-dimensional NODEs using graph-based verification tools.
@@
Other techniques to improve NODE robustness include using skew-symmetric ODE blocks [71] and integrating Guassian process [72], both of which empirically enhance robustness. Though not explicitly designed for adversarial robustness, some efforts contribute similarly by regularizing NODE dynamics. Kinetic and Jacobian norm regularizers [73], along with higher-order derivative regularizers [74], have been proposed to enforce smoothness in NODE dynamics and have been shown to be effective. The method proposed in STEER [75] mitigates the learning of stiff ODEs by randomizing the terminal times during training.
@@
In contrast, there are only handful of works that have investigated security and privacy aspects of NSDEs. A few notable works include Jia et al. [76], who demonstrated that attributions generated from NSDEs are less noisy, visually sharper, and quantitatively more robust than those computed from deterministic NODEs. Liu et al. [32] also showed formally and empirically that NSDEs improve robustness against adversarial examples. More broadly, NSDEs have been widely adopted as a tool to estimate uncertainty [77] and generative modeling [30].
7 Conclusion
@@
This work studies the privacy implications of emerging neural networks, specifically NODEs, through the lens of membership inference attacks. To our knowledge, this is the first work that conducts a comprehensive analysis of the membership risks associated with NODEs. Our key findings are: (1) NODEs are twice as resilient to existing membership inference attacks compared to conventional feedforward models like ResNets, while maintaining comparable accuracy. (2) This privacy benefit stems from reduced model overfitting. NODEs exhibit a smaller generalization gap compared to conventional models, which we attribute to their constrained expressivity—restricting the modeling of underlying dynamics to a system of differential equations. (3) Moreover, advanced, non-stochastic variants of NODEs can also reduce overfitting, leading to both improved accuracy and a further reduction of membership inference risks.
@@
Our analysis does not imply that NODEs are free from membership inference risks. To further mitigate these risks, we study NSDEs, a stochastic extension of NODEs that incorporates an additional diffusion term into system modeling. Most importantly, we formally show that this diffusion term functions as a differentially-private (DP) mechanism, thereby establishing NSDEs as DP-learners. Our theoretical analysis shows that the privacy guarantee ($\varepsilon$) offered by NSDEs is equivalent to that of DP-SGD when using identical mini-batch training configurations. We demonstrate empirically that NSDEs are more effective at reducing membership risks compared to existing defenses that lack provable privacy guarantees. Unlike DP-SGD, NSDEs achieve this protection with a smaller impact on model utility.
@@
Moreover, we present an effective strategy, replace-then-finetune, for using NSDEs as a drop-in privacy-enhancing module. We replace the last few layers of a pre-trained conventional feedforward model with an NSDE block and finetune the adapted model for a few epochs. In evaluation, this strategy reduces the membership inference risks of the original pre-trained model by an order of magnitude, while achieving higher model utility than DP-SGD and requiring only half the training cost (see Appendix for comparisons).
@@
Our work demonstrates that systems learned by neural networks can be effectively modeled with privacy. Our findings imply that advanced neural networks incorporating diffusion mechanisms, such as stable diffusion models, may, with careful modeling, serve as inherently private models while preserving its strong performance. As future work, investigating these private systems may offer valuable insights into how DP shapes what learned by neural networks. We also hope our findings inspire the broader adoption of the replace-then-finetune strategy as a practical and effective method for enhancing model privacy.
Acknowledgments
@@
A. Model Architectures
We adopt architectures examined in prior work [43].
ResNet-14 is defined as
$\begin{split}[&\text{Conv2d}\rightarrow\text{BN}\rightarrow[\text{ResBlock}\times 3\rightarrow\text{BN}\rightarrow\text{Conv2d}]\times 2\\ &\rightarrow\text{ResBlock}\times 3\rightarrow\text{AdaptiveAvgPool2d}\rightarrow\text{Linear}],\end{split}$
@@ A. Model Architectures
Both models begin with Conv2d that processes a 3-channel input image to produce 32 feature maps using a 3$\times$3 filter with a stride of 2 and padding of 1. Following the initial convolutional layer, the features pass through a group of either three ResBlocks or a single ODEBlock, depending on the configuration. Subsequently, the network transitions through a Downsampling layer, which effectively reduces the spatial dimensions while adjust the channel depth, preparing the features for subsequent layers. After the third group of three ResBlocks or an ODEBlock, the network utilizes an adaptive average pooling layer to condenses the feature maps into a single dimensional vector per feature map. This vector is then flattened and passed through a linear layer which acts as the classifier of the network, outputting the final predictions across the categories.
@@
TABLE V: Membership inference risks of ODENet-64 models and SDENet $\sigma\!=\!0$ models. We evaluate ODENet-64 models and and SDENet $\sigma\!=\!0$ models against six existing attacks on CIFAR-10 and CIFAR-100.
@@
Acc. (C-10) Acc. (C-100) TPR @ 0.1% FPR TPR @ 1% FPR AUC Inference acc. Model Method Train Test Train Test C-10 C-100 C-10 C-100 C-10 C-100 C-10 C-100 ODENet-64 Yeom et al. [3] 94.14% 84.36% 70.93% 51.85% 0.00% 0.03% 0.28% 1.17% 0.536 0.659 53.57% 66.09% Shokri et al. [4] 0.08% 0.09% 0.75% 0.85% 0.473 0.480 50.03% 50.01% Song and Mittal [38] 0.01% 0.13% 0.53% 1.58% 0.511 0.497 51.17% 51.63% Watson et al. [34] 0.00% 0.03% 0.28% 1.17% 0.536 0.659 53.57% 66.09% Carlini et al. [6] 0.86% 1.72% 4.17% 7.70% 0.639 0.741 59.57% 69.49% Zarifzadeh et al. [35] 2.72% 2.91% 8.40% 11.47% 0.617 0.757 59.96% 70.65% SDENet $\sigma\!=\!0$ Yeom et al. [3] 89.09% 83.93% 68.01% 52.49% 0.00% 0.09% 0.39% 1.12% 0.538 0.625 54.00% 60.52% Shokri et al. [4] 0.07% 0.05% 0.68% 0.63% 0.468 0.449 50.02% 50.00% Song and Mittal [38] 0.00% 0.14% 0.50% 1.40% 0.516 0.496 51.65% 51.13% Watson et al. [34] 0.00% 0.09% 0.44% 1.12% 0.538 0.625 54.00% 60.52% Carlini et al. [6] 0.35% 1.03% 2.02% 5.36% 0.559 0.676 54.09% 62.26% Zarifzadeh et al. [35] 0.91% 2.04% 4.01% 8.04% 0.579 0.701 55.76% 63.58%
B. Differential Privacy of Gaussian Mechanism